40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 29 03:16:21 2016, ericq, Update, LSC, DRFPMI Locked Once More 2016-03-29_calibdarm.pdf2016-03-29_DRFPMI_errctrl.pdfDRFPMI_DTT.zip
    Reply  Wed Mar 30 16:40:24 2016, ericq, Update, LSC, 2016 vs 2010 2016_vs_2010.png
    Reply  Mon Apr 11 16:06:30 2016, ericq, Update, LSC, DRFPMI Data Archived 
Message ID: 12055     Entry time: Wed Mar 30 16:40:24 2016     In reply to: 12053
Author: ericq 
Type: Update 
Category: LSC 
Subject: 2016 vs 2010 

I haven't found any data files for the DARM spectrum of the previous generation of 40m, but with some GIMP-fu, I have plotted Monday's spectrum (green) on top of one of the figures from Rob's thesis.

ELOG V3.1.3-